# Negation: A Pink Elephant in the Large Language Models’ Room?

### v1

@@ Abstract
We construct two multilingual natural language inference (NLI) datasets with paired examples differing in negation. We investigate how model size and language impact its ability to handle negation correctly by evaluating popular LLMs.
Contrary to previous work, we show that increasing the model size consistently improves the models’ ability to handle negations. Furthermore, we find that both the models’ reasoning accuracy and robustness to negation are language-dependent and that the length and explicitness of the premise have a greater impact on robustness than language.
Our datasets can facilitate further research and improvements of language model reasoning in multilingual settings.
@@ 1 Introduction
The negation poses a fundamental problem for large language models, as shown by many studies Kassner and Schütze (2020); Truong et al. (2023a); García-Ferrero et al. (2023); Nadeem et al. (2024); Kalouli et al. (2022); Zhang et al. (2023). The phenomenon is called “NO syndrome” or “negation blindness”.
Vision-Language models do not understand negation either Alhamoud et al. (2025): when asked to provide a picture with no elephant in the room, they put an elephant there anyway. It appears that negation tokens (e.g., “not”) have a limited effect on the representations learned distributionally.
However, in tasks requiring logical reasoning, processing negation is critical for reliable performance.
@@
In this paper, we evaluate LLMs on the text entailment (TE) task, exploring the ability of current LLMs to reason under negation. As in the known story about the elephant met by six blind monks, the problem of LLMs dealing with negation has been touched on from a number of different angles and points of view. Hermann et al. (2013) pioneeringly sees shortcomings in the ability of language models to capture logical operations such as negation in continuous vector space representation of compositional distributional semantics. Other points of view are the inclusion of more negative training data Helwe et al. (2022), modification of the input text itself to remove negation whilst keeping the overall meaning of the text Rezaei and Blanco (2024), or the fine-tuning of the instruction prompt to explicitly include the possible presence of negation in the texts Truong et al. (2023a). Ettinger (2020) assessed BERT’s Devlin et al. (2019) sensitivity to the meaning of negation by measuring the proportion of items in which the model assigns higher probabilities to true completions than to false ones. Kassner and Schütze (2020) studied negation and mispriming effects for pre-trained language models. Models memorize both positive and negative facts but fail to generalize to unseen facts. Dziri et al. (2023) investigate limits of LLMs across three representative compositional tasks.
In this paper, on the other hand, in the series of experiments, we pin down and separate the primary causes of the problem:
@@
In one of the findings of Truong et al. (2023a), it is posited that the LLMs are unable to reason under negation, with the increased size of the LLM not being a mitigating factor. We review this claim by evaluating models from the same LLM family with different model sizes.
@@
Can the adjustments in architecture and model refinements improve models’ capabilities to handle negations?
Furthermore, we evaluate whether the difference in architecture is a deciding factor in the model’s ability to understand negation.
RQ3:
@@
2 Methodology
We base our evaluation on the FEVER Thorne et al. (2018a) and SNLI Bowman et al. (2015) datasets. From these English NLI datasets, we prepared our variants of two datasets in four languages through the pre-processing and translation processes. We have extended the original English datasets with negated versions of the hypotheses, using our negation pipeline. We have manually reviewed the correctness of these hypotheses, manually modifying the sentences in cases where the pipeline fell short. We then translated these English datasets using DeepL and validated the quality of translation by manually checking 100 random samples from each dataset. Furthermore, we also evaluated a preexisting Czech translation of the FEVER-NLI dataset, which has been fully verified by a human annotator; we denote this dataset in our results as CES FEVER MANUAL. More details of dataset preparation are described in Appendix A.
All of the models evaluated are publicly available state-of-the-arg models based on the transformer architecture Vaswani et al. (2017), have open weights, a multi-lingual pre-training, and are fine-tuned to follow user instructions. In the evaluation, we prompt the model to judge if a hypothesis follows from a premise or not. We let the model finish a response “The answer is:” with either True or False using a greedy decoding. To ensure the response follows the predefined structure, we employ Outlines Willard and Louf (2023) and LMFE Gat (2024). More details of the prompt format can be found in Appendix C.
3 Results
We summarize our results in this section. We find that accuracy and negation sensitivity improve with increasing model size and the information specificity of dataset premises, while variations in model architectures and languages are not statistically significant (for details, see Appendix D.
Figure 4 shows the accuracy change of different models for the hypotheses with and without negation. We found that all evaluated models perform better than a random guess and that the size of the model has a strong positive correlation with the model’s ability to correctly interpret negation. The larger models narrow the gap between the accuracies, with two outliers on the SNLI dataset: a slight increase on the middle-sized Llama model (8B) and a marked drop on the middle-sized Mistral model (22B).
The difference in our findings can be caused by a particular choice of the evaluated models. However, a plausible explanation is that measured accuracy strongly benefits from constraining output to a set of predefined options rather than expecting the language model to adhere by itself to a desired output format.
The overall findings are in contrast to the Finding of Truong et al. (2023b), where it is posited that pre-trained language models perform comparably or worse than a random baseline, and model scaling has almost no effect.
Larger LMs are more sensitive to negation for TE task.
As we can see in Table 5 in the Appendix, the size of the model has a positive correlation with the overall negation sensitivity of the models. Larger models are less likely to classify pair hypotheses with the same textual entailment label, and the impact of negation is more effectively propagated through the model. One outlier in our results is the Mistral (22B) model, which has outperformed the Mistral (12B) model on CES, ENG, and UKR SNLI datasets.
Architecture adjustments do not substantially impact models’ ability to handle negation.
Apart from differences in size, performance across different LLM families are comparable.
This suggests that, on average, current architectural differences in state-of-the-art models may not be the deciding factor in a model’s ability to understand negation. Instead, it seems that the size of the model plays a prominent role in determining how well it handles negations, likely due to its increased capacity to capture nuanced language patterns. In Appendix D.4, we further discuss the influence of model size on the models’ ability to more accurately model the language patterns.
Models’ accuracy is not language-dependent.
Figure 4 shows the accuracy change on different models with regard to the language of the dataset. Our results do not fully conform to our initial presuppositions that the models will perform worse in languages with higher word-order flexibility. The Czech datasets have generally performed worst out of the evaluated datasets, with the CES FEVER MANUAL achieving higher accuracy on hypotheses without negation than the CES FEVER dataset.
However, in conjunction with Figures 4 and Figure 4, our results lead us to conclude that the length and specificity of the premise have a larger impact than the language of the evaluated dataset. Whereas the models were able to properly handle negation on the FEVER dataset, which is based on a fact verification dataset and thus has rich and factually specific premises, the negation on the SNLI dataset, which is more dependent on common sense knowledge and implicit semantical patterns the models learn during training, has shown itself to be problematic, even for large models.
4 Conclusions and future work
This work presents a novel dataset enabling the evaluation of language models’ robustness in handling negations in four languages.We have extended an existing FEVER Thorne et al. (2018b) and SNLI Bowman et al. (2015) datasets to include paired hypotheses that are logically opposite to each other by negating the primary verb in the original hypothesis. We have evaluated the models’ accuracy separately on hypotheses without negation and hypotheses with negation computed the accuracy drop that occurred due to the addition of negation in the hypothesis.
We find that model size plays a key role in handling negation correctly, countering the conclusions of previous work Truong et al. (2023a). However, there seems to be minimal effect on the accuracy that can be attributed to the different architectures and pre-training data. The results on differing language datasets only partially match the initial expectations, with models performing worst on the Czech configuration; however, datasets created using machine translation performed similarly, regardless of the language morphology. The accuracy more heavily depended on the specificity of information in the premise; the more common sense SNLI dataset performed worse than the more fact-explicit FEVER dataset.
Our work provides new resources for addressing the deficiency of language models in handling negations. Our methodology will be used to find the root cause of the negation problem by computing the accuracy on subsets of our new datasets in 4 languages (e.g., a subset of nonprojective expression of negation, a subset of different ways of expressing syntactic or semantic negation Kalouli et al. (2022) or to evaluate sub-clausal negation based on a dataset of Truong et al. (2022). We will use the lense of LLM-microscope Razzhigaev et al. (2025) to look on the tokens that play decisive role.
Limitations
The details of model preparation (training data, hyperparameters, fine-tuning) limit the full understanding of the results we report. Also, the results might be biased by the [semi]automated translations of datasets to their German, Czech, and Ukrainian versions. However, we did a manual check of the quality of a subset of the translations and did not encounter any errors related to the application of negation in context.
Appendix A Dataset preparations
@@
We evaluate the accuracy of the models in our experiments using two datasets formatted for the NLI task, i. e. containing premise-hypothesis pairs and the label denoting their logical relationship (entailment, neutral, contradiction):
@@
Thorne et al. (2018a) is a modification of the FEVER dataset, which originally focused on the task of fact extraction and verification. The dataset consists primarily of factual statements retrieved from Wikipedia.
As our aim is to study the effects of negation on the accuracy of the models, we extend these datasets to contain the negation of the original hypothesis. The original datasets contain a low percentage of hypotheses with common negation markers, such as no, n’t, not, nobody, neither, or nothing, and differ in the verbosity of premises, as seen in the Table 1.
In Table 2, we include average lengths of premises and hypotheses of the newly prepared datasets for comparison.
@@
Table 2: Comparison of average lengths across evaluated datasets
A.1 Preparation of new datasets in English
In order to prepare new datasets containing negation, we first sanitize the datasets, and then we extend the datasets with hypotheses, including negation. We then manually replace incorrectly generated sentences and create two datasets: the EngNoFEVER dataset, containing 2,600 premise-hypothesis-negated_hypothesis triplets, and the EngNoSNLI dataset, containing 6,261 triplets.
@@ A.1 Preparation of new datasets in English
Firstly, we sanitize the datasets by removing all incomplete or invalid rows, i.e., removing rows with empty premises, empty hypotheses, or invalid labels. Secondly, we remove rows that have the neutral entailment label as the effect of negation on their entailment value is not uniform.
@@
We employ the negate Python module for our negation. This module is a rule-based negation tool for the English language, utilizing the spaCy model en_core_web_md for POS tagging and dependency parsing in order to correctly locate the verb in the sentence. The verb is then negated by the inclusion of an auxiliary verb and a negation marker, with our preference in the module set to the contraction n’t. The tool is able to negate in both directions, meaning that it can remove negation from the verb as well.
@@
As the tool we used for the hypothesis generation is rule-based, it comes with certain limitations and cases that are not covered by the rules and have to be fixed manually.
@@
The rules of the module as aimed the Not-negation, most likely due to its higher frequency in the English language, in comparison to the No-negation.
@@
A.2 Translation of new datasets into Czech, German, and Ukrainian
To explore the impact of negation on the accuracy of the LLMs across languages, we translate our English (ENG) datasets into Czech (CES), German (DEU), and Ukrainian (UKR). We have chosen these languages due to their differing features, such as morphology, orthography, alphabets, and projectivity, as specified in Table 3.
@@ A.2 Translation of new datasets into Czech, German, and Ukrainian
After we had translated the datasets into our target languages, we randomly selected 100 rows from each dataset and checked the quality of the translation.
Although the translation quality of the model we have used is generally of high quality, it is still an LLM that is not immune to hallucination on rarely seen data. We have, for example, in the Ukrainian SNLI dataset, observed the outlier case shown in Table 4. This behavior does not occur when the sentence does not include negation. We suppose that the combination of the inclusion of negation, the tense in the original English sentence, and the inclusion of an initialism (pb and j) created confusion in the model. However, as the training data of the model isn’t public, we cannot confirm our hypothesis.
@@
Appendix B Full details of accuracy across evaluated models
We include the full tables of accuracies of models for hypotheses with and without negation on the task of textual entailment. In both tables, we can see that almost all models achieve the highest accuracy on the English datasets, on both hypotheses, with and without negation.

### v2

@@ Abstract
We constructed and published two new textual entailment datasets NoFEVER-ML and NoSNLI-ML in four languages (English, Czech, German, and Ukrainian) with paired examples differing in negation. It allows investigation of the root causes of the negation problem and its exemplification: how popular LLM model properties and language impact their inability to handle negation correctly.
Contrary to previous work, we show that increasing the model size may improve the models’ ability to handle negations. Furthermore, we find that both the models’ reasoning accuracy and robustness to negation are language-dependent and that the length and explicitness of the premise have an impact on robustness.
There is better accuracy in projective language with fixed order, such as English, than in non-projective ones, such as German or Czech.
Our entailment datasets pave the way to further research for explanation and exemplification of the negation problem, minimization of LLM hallucinations, and improvement of LLM reasoning in multilingual settings.
@@ 1 Introduction
Negation, a cornerstone of logical reasoning and nuanced communication, often remains a ‘pink elephant in the room’ for Large Language Models (LLMs). Despite their impressive advancements, effectively understanding and processing negation poses a fundamental, and often underestimated, challenge to their reliability and logical reasoning capabilities. Processing negation is critical for reliable performance, and leads to hallucinations as Varshney et al. (2025) showed on four tasks with negation: ‘false premise completion’, ‘constrained fact generation’, ‘multiple choice question answering’, and ‘fact generation’. Vision-language models do not understand negation either Alhamoud et al. (2025): when asked to provide a picture with no elephant in the room, they put an elephant there anyway. It appears that negation tokens (e.g., “not”) have a limited effect on the representations learned distributionally Singh et al. (2024).
This difficulty, sometimes termed negation blindness or NO syndrome, is well-documented. Studies show LLMs may struggle to distinguish between facts and their negations, misunderstand the semantic impact of negative particles, and fail to generalize negation handling robustly, even with instruction tuning.
@@
Prior Work on Negation in LLMs
The challenge of negation for LLMs has been approached from several angles, as in the parable about the elephant met by blind monks. We categorize prior work into the following areas:
1. Foundational Issues & Early Observations
Early research highlighted fundamental difficulties for language models in capturing logical operations such as negation within continuous vector space representations Hermann et al. (2013). Subsequent work observed that LLMs can be equally prone to generating factual statements and their incorrect negations Kassner and Schütze (2020), and specific assessments were made on models like BERT regarding their sensitivity to the meaning conveyed by negation Ettinger (2020).
2. Datasets & Benchmarks for Negation
Several benchmarks have been developed to probe negation understanding. For instance, the NeQA dataset was introduced to test question answering with negation, where models did not exhibit straightforward positive scaling Zhang et al. (2023). Another large-scale benchmark was created to challenge LLMs with a wide array of negative sentences, finding that while fine-tuning helps, generalization in handling negation remains elusive García-Ferrero et al. (2023). The NaN-NLI test suite specifically focuses on sub-clausal negation Truong et al. (2023). Mondorf and Plank (2024) introduced a benchmark for suppositional reasoning that convincingly shows that “proficient models struggle with inferring the logical implications of potentially false statements”.
3. Analysis of LLM Failures & Characteristics of Negation Handling
Researchers have investigated whether the semantic constraints of function words like negation are adequately learned and how context impacts their embeddings Kalouli et al. (2022). It has been shown that LLMs tend to underestimate the significant impact of negation on sentence meaning Anschütz et al. (2023a). A broad analysis across various LLMs, including instruction-tuned models, confirmed their inability to capture the lexical semantics of negation and to reason effectively under negation, with model size not initially appearing as a clear mitigating factor Truong et al. (2023) (this paper re-evaluates the model size aspect). Furthermore, negation tokens like “not” appear to have a limited effect on the distributional representations learned by models Singh et al. (2024). The phenomenon often leads to hallucinations in tasks involving negation, such as false premise completion and constrained fact generation Varshney et al. (2025). Other studies have also investigated the limits of LLMs on general compositional tasks, which implicitly include negation Dziri et al. (2023).
4. Negation in Multimodal & Vision-Language Models
The challenge of negation extends beyond text-only models. Vision-language models have also been shown to struggle with understanding negation, for example, when prompted to generate images without a certain object Alhamoud et al. (2025). Efforts are being made to address this, such as proposing negation context-aware reinforcement learning feedback loops to ensure generated text and images correctly reflect negated queries Nadeem et al. (2024).
5. Mitigation Strategies & Improving Negation Handling
Various strategies have been proposed to improve negation handling. These include augmenting training data with more negative examples Helwe et al. (2022), paraphrasing input text into affirmative terms to remove explicit negation while preserving meaning Rezaei and Blanco (2024), and exploring fine-tuning processes, including instruction prompt adjustments, to explicitly account for negation Anschütz et al. (2023b); Truong et al. (2023). Castricato et al. (2024) study controllable generation of LLM through the lens of the “Pink elephant paradox” Spiers (2002) and think that form of Reinforcement Learning from AI Feedback (RLAIF) that they term Direct Principle Feedback assesses the problem.
6. Cross-lingual & Multilingual Negation Studies
The study of negation awareness has also been extended to multilingual contexts. For example, Hartmann et al. (2021) developed a multilingual benchmark to probe negation-awareness across languages including English, Bulgarian, German, French, and Chinese, setting a precedent for cross-lingual investigations in this domain. Vrabcová and Sojka (2024) develop CsFEVER dataset and confirm that negation causes problems in Czech as well.
In this paper, we focus on the effects of moderating variables on robustness to negation, posed by the following research questions:
@@ 6. Cross-lingual & Multilingual Negation Studies
In one of the findings of Truong et al. (2023), it is posited that the LLMs are unable to reason under negation, with the increased size of the LLM not being a mitigating factor. We review this claim by evaluating models from the same LLM family with different model sizes.
@@
The primary contributions of this work are:
  • The development and release of two novel, multilingual textual entailment datasets specifically designed to evaluate negation. These datasets span English, Czech, German, and Ukrainian and feature paired examples differing only in negation.
  • Empirical evidence demonstrates that larger model sizes can indeed improve the handling of negations, offering a counterpoint to some earlier research. An analysis reveals that while linguistic features of a language play a role, the specificity and length of the premise context have a more pronounced impact on an LLM’s robustness to negation than the language itself.
The paper is structured as follows. Section 2 presents the preparation of our two new NLI datasets. Section 3 describes the methodology and the LLMs we use in our experiments. Results are discussed in Section 4. We conclude and outline future directions in Section 5.
2 Datasets
@@
We evaluate the accuracy of the models in our experiments using two datasets that have been originally formatted for the Natural Language Inference (NLI) task, i.e., containing premise-hypothesis pairs and the label denoting their logical relationship (entailment, neutral, contradiction).
@@
Thorne et al. (2018) is a modification of the FEVER dataset, which originally focused on the task of fact extraction and verification. The dataset consists primarily of factual statements retrieved from Wikipedia.
As our aim is to study the effects of negation on the accuracy of the models, we modify the format of these datasets for the purpose of the TE task and extend them to contain the negation of the original hypothesis. The original datasets contain a low percentage of hypotheses with common negation markers, such as no, n’t, not, nobody, neither, or nothing, and differ in the verbosity of premises, as seen in the Table 1.
In Table 2, we include average lengths of premises and hypotheses of the newly prepared datasets for comparison, NoFEVER-ML and NoSNLI-ML.
@@
Table 2: Comparison of average lengths across evaluated datasets, NoFEVER-ML and NoSNLI-ML.
2.1 Preparation of new datasets in English
In order to prepare new datasets containing negation, we first sanitize the datasets and then extend the datasets with hypotheses including negation. Contrary to Gururangan et al. (2018), we do not distinguish between contradiction and logical negation of tokens expressing negation. We then manually replace incorrectly generated sentences and create 2 datasets.
NoFEVER-ENG dataset:
This dataset contains 2,600 premise-hypothesis-negated_hypothesis triplets, with 1,513 premises entailing the hypothesis without negation, and 1,087 premises entailing the hypothesis with negation.
NoSNLI-ENG dataset:
This dataset contains 6,261 triplets, with 3,245 premises entailing the hypothesis without negation, and 3,016 premises entailing the hypothesis with negation.
@@ 2.1 Preparation of new datasets in English
Firstly, we sanitize the datasets by removing all incomplete or invalid rows, i.e., removing rows with empty premises, empty hypotheses, or invalid labels. Secondly, we remove rows that have the neutral entailment label, as the effect of negation on their entailment value is not uniform.
@@
We employ the negate Python module for the first phase of our negation process. This module is a rule-based negation tool for the English language, utilizing the spaCy model en_core_web_md for POS tagging and dependency parsing in order to correctly locate the verb in the sentence. The verb is then negated by the inclusion of an auxiliary verb and a negation marker, with our preference in the module set to the contraction n’t. The tool is able to negate in both directions, meaning that it can remove negation from the verb as well.
@@
As the tool we used for the hypothesis generation is rule-based, it comes with certain limitations and cases that are not covered by the rules and have to be fixed manually. For the second phase of our negation process, we have performed a full manual check to ensure the correctness of the newly created hypotheses.
@@
The rules of the module as aimed at the Not-negation, most likely due to its higher frequency in the English language, in comparison to the No-negation.
@@
2.2 Translation of new datasets into Czech, German, and Ukrainian
To explore the impact of negation on the accuracy of the LLMs across languages as started by Hartmann et al. (2021) for English, Bulgarian, German, French and Chinese, we translate our English (ENG) datasets into Czech (CES), German (DEU), and Ukrainian (UKR). We have chosen these languages due to their differing features, such as morphology, orthography, alphabet, and projectivity, as specified in Table 3.
@@ 2.2 Translation of new datasets into Czech, German, and Ukrainian
We carry out the Statistical Sign Test with the ties equally split and with the alpha level $\alpha=0.05$. The null hypothesis is that both translators are of equal quality. Our result value is $p$-value = $0.2$ and thus we cannot reject our null hypothesis.
@@
After we had translated the datasets into our target languages, we randomly selected 600 rows from each dataset and checked the quality of the translation.
Although the translation quality of the model we have used is generally high, it is still an LLM that is not immune to hallucinations on rarely seen data. We have, for example, in the Ukrainian SNLI dataset, observed the outlier case shown in Table 4. This behavior does not occur when the sentence does not include negation. We suppose that the combination of the inclusion of negation, the tense in the original English sentence, and the inclusion of an initialism (pb and j) created confusion in the model. However, as the training data of the model isn’t public, we cannot confirm our hypothesis.
Dataset Availability
Our final datasets are publicly available in an anonymized repository Anonymous (2025).
@@
3 Methodology
All of the evaluated models are publicly available state-of-the-art models based on the transformer architecture Vaswani et al. (2017), have open weights, a multi-lingual pre-training, and are fine-tuned to follow user instructions. In the evaluation, we prompt the model to judge if a hypothesis follows from a premise or not. We let the model finish a response “The answer is:” with either True or False using a greedy decoding. To ensure the response follows the predefined structure, we employ Outlines Willard and Louf (2023) and LMFE Gat (2024). More details of the prompt format can be found in Appendix B.
4 Results
We summarize our results in this section, including the results of the statistical tests to support these claims. We find that accuracy and negation sensitivity improve with increasing model size and the information specificity of dataset premises, while variations in languages are not statistically significant.
Unless otherwise stated, for the statistical tests, we are using an alpha level (significance threshold) $\alpha=0.05$. More details of the specific values used during computation can be found in Appendix C.
Figure 2 shows the accuracy change of different models for the hypotheses with and without negation. We found that all evaluated models perform better than a random guess and that the larger models narrow the gap between the accuracies, with two outliers on the SNLI dataset: a slight increase on the middle-sized Llama model (8B) and a marked drop on the middle-sized Mistral model (22B).
Using the results shown in Figure 4, we compute the correlation of the model size and the relative accuracy change for each evaluated dataset, as well as the overall correlation. Using the Shapiro-Wilk test, we find that the majority of datasets do not have a normal distribution over the different model sizes. Thus, we compute correlation using the Spearman correlation coefficient, as the Pearson correlation coefficient presumes a normal distribution of values.
The results indicate a strong positive correlation between model size and the relative accuracy change on the FEVER dataset and a weak-to-moderate positive correlation on the SNLI dataset. For an averaged relative accuracy change for all languages per model, the results indicate a strong positive correlation with the Spearman correlation coefficient of $0.867$, supporting out RQ1.
Our findings are in contrast to the results of (Truong et al., 2023, Section 3, Finding 1), where it is posited that pre-trained language models perform comparably or worse than a random baseline, and model scaling has almost no effect. The difference in our findings can be caused by a particular choice of the evaluated models and their level of fine-tuning. A plausible explanation might be that measured accuracy strongly benefits from fine-tuning of prompts.
Larger LMs are more sensitive to negation for TE task.
Table 5 shows the negation sensitivity of each model across individual datasets. Our negation sensitivity metric represents the percentage of datasets in which the model assigned opposite entailment scores to pairs of hypotheses with and without negation. The best results were achieved only by the biggest models, supporting our RQ1.
Larger models are less likely to classify pair hypotheses with the same textual entailment label, and the impact of negation is more effectively propagated through the model. One outlier in our results is the Mistral (22B) model, which has been outperformed by the Mistral (12B) model on the CES, ENG, and UKR SNLI datasets, as well as on the DEU FEVER dataset.